Loreto Inc. has the following financial ratios: asset turnover = 2.00; net profit margin (i.e., net income/sales) = 4%; payout ratio = 35%; equity/assets = 0.30.

a. What is Loreto's sustainable growth rate?

b. What is its internal growth rate?

(Do not round intermediate calculations. Enter your answers as a percent rounded to 2 decimal places.)
Answer #1

Sustainable Growth rate = Retention ratio *Return on Equity /(1-(Retention ratio *Return on Equity))

From Dupont

Return on Equity = Profit Margin * Asset Turnover * Leverage ratio

Where profit Margin = 4%, Asset Turnover = 2, Leverage ratio = Assets/ equity = 1/Equity/Assets = 1/0.30 = 3.3333

Return on Equity = 4%*2*3.333 = 26.6667%

Retention ratio = 1-Payout ratio = 1-35% = 0.65

Sustainavle Growth rate = (0.65*0.266667)/(1-(0.65*0.266667)) = 0.1733/0.8267 = 0.209677

Therefore Sustainable Growth rate = 20.97%

Internal Growth rate = (Retention ratio * Return on Assets)/(1-(Retention ratio * Return on Assets))

Return on Assets = Profit / Assets

Given Asset Turnover = Sales/Assets = 2

Profit / Sales = 4%

By multiplying both

(Sales / Assets) *(Profit/ Sales) = 2*4%

Profit / Assets = 8%

Internal Growth rate = (0.65*0.08)/(1-(0.65*0.08) = 0.052/0.948 = 0.05485

Therefore Internal Growth Rate = 5.49%